Spiderman makes a leap from one building to another. He starts on one building that is 68 m tall and lands on a building that is 54 m tall. It takes him 0.9 s to make the jump. The gap between the buildings is 18 m.


What is the magnitude and direction of his velocity as he jumps?


What is the magnitude and direction of his velocity as he lands?
Answer #1

Horizontal velocity Ux = 18/0.9 = 20 m/s

using second equation of motion in vertical, h = Uyt - 0.5gt^2

54-68 = Uy*0.9 - 0.5*9.8*0.9^2

Uy = -11.15 m/s

Final vertical velocity = Uy - gt = -11.15 - 0.9*9.8 = - 19.97 m/s

Jumping speed u = sqrt(20^2+11.15^2)

=   22.9 m/s

direction = arctan(11.5/20) = 29.9 degree below horizontal

landing speed v = sqrt(20^2+19.97^2)

= 28.2 m/s

direction = actan(19.97/20) = 45 degree below horizontal